# Variability of the UV luminosity function with SPICE

### v1

@@ Abstract
pubyear: 2015pagerange: Variability of the UV luminosity function with SPICEB
@@ 1 Introduction
Since the James Webb Space Telescope (JWST) started to provide revolutionary data in July 2022, our understanding of the early galaxy formation process has been challenged. Although over the past decade the Hubble Space Telescope (HST) enabled us to observe galaxies out to $z\approx 10$ (Zheng et al., 2012; Coe et al., 2013; Oesch et al., 2016; Morishita et al., 2018; Bagley et al., 2022), JWST has pushed our observational barrier to even higher redshifts, and has detected an unexpected overabundance of massive galaxies at $z>10$ (Castellano et al., 2022; Finkelstein et al., 2022; Naidu et al., 2022; Adams et al., 2023; Morishita & Stiavelli, 2023; Bouwens et al., 2023a; Bouwens et al., 2023b; Donnan et al., 2023a; Atek et al., 2023; Pérez-González et al., 2023; Willott et al., 2024; Donnan et al., 2024). This discovery might challenge theoretical models, as it suggests a high incidence of bright galaxies at this epoch (Mason et al., 2023; Boylan-Kolchin, 2023; Lovell et al., 2023; Acharya et al., 2024). The recent detection of a spectroscopically confirmed galaxy at $z=14.32$ by Helton et al. (2024) significantly elevates the challenge, suggesting that such bright objects exist even at these high redshifts. These observations raise several intriguing questions: Could there be a missing population of bright stars? Is there something unique about the early star formation history? Do these findings create tensions with the standard $\Lambda$CDM cosmology?
@@ 2 The SPICE simulation suite
SPICE has been performed with RAMSES-RT (Rosdahl et al., 2013; Rosdahl & Teyssier, 2015), the radiation-hydrodynamics extension of the Eulerian Adaptive Mesh Refinement (AMR) code RAMSES (Teyssier, 2002). The simulations target a cubic box of size 10 $\rm{\mathit{h}^{-1}cMpc}$ with $512^{3}$ dark matter particles of mean mass $m\rm{{}_{dm}}=6.38\times 10^{5}\mathrm{M_{\odot}}$. The AMR allows each parent gas cell to split into 8 smaller cells when certain conditions are met (see B24). As a reference, this provides a physical resolution at $z=5$ ranging from 4.8 kpc for the coarsest level ($\ell_{\text{min}}=9$) to approximately 28 pc for the finest level ($\ell_{\text{max}}=16$). The hydrodynamical equations are solved using a second order Godonov scheme and the dynamics of collisionless dark matter and stellar particles are computed by solving the Poisson equation with a particle-mesh solver, projecting onto a grid using the Cloud-in-Cell scheme (Guillet & Teyssier, 2011). Gas cooling and heating are accounted for as in Rosdahl et al. (2013). Star-formation has been modelled following Kretschmer & Teyssier (2020), which adopt a subgrid turbulence model depending on the physical conditions of each computational cell including parameters like the local gas turbulent Mach number and the virial parameter, which acts as a measure of the gravitational balance in the region, and which results in a spatially-variable star formation efficiency. Additionally, SPICE includes a radiative transfer scheme with five frequency groups, i.e. infrared (0.1-1 $\rm{eV}$), optical (1-13.6 $\rm{eV}$), and three UV (13.6-24.59 $\rm{eV}$; 24.59-54.42 $\rm{eV}$; 54.42$\infty$ $\rm{eV}$).
@@
  • "smooth-sn": for a given stellar population, a SN with mass $>20\rm{M_{\odot}}$ can explode as early as 3 Myr, whereas SN with progenitor mass of $8\rm{M_{\odot}}$ can explode as late as 40 Myr. An energy of $2\times 10^{51}$ ergs per SN explosion is injected into the neighbouring cells.
@@ 3.1 UV luminosity function
Figure 1: 1500 $\mathrm{\SIUnitSymbolAngstrom}$ luminosity functions at $z$ = 12 (top panel) and 11 (bottom panel) for three feedback models (in different colors). Solid and dashed curves refer to intrinsic and dust attenuated LFs, respectively. A compilation of observations from HST and JWST (Bouwens et al., 2015; Harikane et al., 2022; Naidu et al., 2022; Adams et al., 2023; Harikane et al., 2023; Bouwens et al., 2023a; Bouwens et al., 2023b; Leung et al., 2023; Donnan et al., 2023b; Donnan et al., 2023c; Pérez-González et al., 2023; Casey et al., 2024; Robertson et al., 2024; McLeod et al., 2024; Whitler et al., 2025) is shown as gray data points.
While in B24 we showed results at $z\leq 10$, in Figure 1 we present the 1500 Å LF for all feedback models at $z=11$ and 12. The LFs have been derived from the stellar spectral energy distribution within a 10 $\mathrm{\SIUnitSymbolAngstrom}$ bin centered around 1500 $\mathrm{\SIUnitSymbolAngstrom}$ We use the SED model of BPASSv2.2.1 (Eldridge et al., 2017; Stanway & Eldridge, 2018), assuming the Chabrier (2003) IMF (details are mentioned in B24). The solid and dashed lines in the figure indicate the intrinsic and dust-attenuated luminosity functions, respectively. To account for dust attenuation, we utilize the Monte Carlo radiative transfer code RASCAS (Michel-Dansac et al., 2020). In this approach, 100 rays are cast from each stellar particle within a halo out to its virial radius. The dust attenuation along each ray is computed following the model described in the previous section.
As already noted in B24, also at these redshifts smooth-sn produces a higher intrinsic luminosity function (by 0.3-0.4 dex) at all magnitudes. Conversely, the hyper-sn model yields the lowest galaxy count across all magnitude bins. The bursty-sn model produces results which are typically in between those of the other models. The effect of dust attenuation (dashed curves) starts to be visible from an absolute magnitude of $M_{1500}\approx-15$, the value changing slightly depending on feedback model and redshift. When comparing the dust attenuated LFs to a compilation of observations from HST and JWST (Bouwens et al., 2015; Harikane et al., 2022; Naidu et al., 2022; Adams et al., 2023; Harikane et al., 2023; Bouwens et al., 2023a; Bouwens et al., 2023b; Leung et al., 2023; Donnan et al., 2023b; Donnan et al., 2023c; Pérez-González et al., 2023; Casey et al., 2024; Robertson et al., 2024; McLeod et al., 2024; Whitler et al., 2025), we notice a fairly good match in the range of magnitudes covered by SPICE (except for the hyper-sn model), which, given the limited box size, does not extend to the brightest observed galaxies. However, to assess whether the extrapolated UVLF aligns with observations, we parametrize and fit the curves using the commonly adopted Schechter (1976) function (see Appendix B). For a more detailed discussion of the LF, we refer the reader to B24.
@@
To investigate the temporal evolution of the dust-attenuated UVLF on short timescales, in Figure 3 we present the UVLF in intervals of 10, 30 and 50 Myr around $z\approx 10$ and 8. While the general behaviour in the three models is similar, we find that the UVLF in the bursty-sn model experiences the most pronounced fluctuations, with a deviation of more than 1 dex. In comparison, the smooth-sn model exhibits a more gradual evolution of the UVLF, producing a steady increase in the number of galaxies. As already mentioned in B24, this suggests a smoother star formation process due to the presence of gentler SN feedback. The hyper-sn model shows an intermediate behaviour at $z\approx 10$, with variations up to 0.5 dex. At $z\approx 8$, the variability in hyper-sn becomes slightly lower than in smooth-sn. In the following sections we will investigate more in detail the origin of such variability.
@@
[…]
@@ 3.2 Star formation rate variability
For a more quantitative investigation of the SFR variability, we first compute the median star-formation main sequence in different halo mass bins and redshift intervals (corresponding to bins of 100 Myr [^1] This has been calculated from the SFR values of all the halos present in each 100 Myr time intervals and then computing the median SFR as a function of halo mass. We have also performed the same analysis using bin sizes of 30 and 50 Myr, and found no qualitative differences in our conclusions.), SFRmedian. In the inset of Figure 4 we show the variability of the SFR for the same halo, defined as $\rm{\mathit{\delta}_{SFR}}={\rm log}_{10}\rm{(SFR/SFR_{median})}$, which quantifies how the SFR of the halo deviates from the median SFR computed over the entire sample of halos in the same mass bin. In the inset of Figure 4 we show the distribution of $\delta_{\rm SFR}$ for this halo during its entire lifetime, together with the corresponding standard deviation, $\sigma_{\rm SFR}$. As expected, bursty-sn shows the widest distribution, with $\sigma_{\rm SFR,bu}=0.34$, which is more than double the one of smooth-sn with $\sigma_{\rm SFR,sm}=0.12$. hyper-sn produces a $\delta_{\rm SFR}$ distribution which is similar to the one from smooth-sn, with a slightly higher standard deviation of $\sigma_{\rm SFR,hn}=0.14$, reflecting the stronger fluctuations present at $z>9$.
@@ 3.3 Variability of the UV luminosity function
Figure 7 shows the scatter in UV luminosity, $\sigma_{\rm UV}^{\rm dust}$, as a function of $M_{\rm halo}$ in the range $z=5-15$. At the highest redshifts, all models exhibit a similar dependence on halo mass, with smaller halos showing a larger UVLF variability, likely due to efficient SN feedback in this mass regime (Gelli et al., 2024). As expected, bursty-sn has the highest variability among all models, with a UV luminosity scatter which starts to rise at $z\lesssim 10$ and reaches $\sigma_{\rm UV}^{\rm dust}\approx 2.5$ for $M_{\rm halo}\approx 10^{10.3}$ M at $z\approx 6.5$. At lower redshift, $\sigma_{\rm UV}^{\rm dust}$ decreases again. In comparison, for smooth-sn the UV luminosity scatter is always below $\sigma_{\rm UV}^{\rm dust}\approx 1.3$, whereas hyper-sn shows slightly higher variability, with a maximum value of $\sigma_{\rm UV}^{\rm dust}\approx 1.5$. Indeed, hyper-sn exhibits a variability comparable to the one in bursty-sn until $z\approx 9$, while at lower redshift the variability decreases as a result of the reduced fraction of HN, and becomes more similar to the smooth-sn one. The same behaviour had already been observed in Figure 4 with respect to the SFR. At $z<6$ all models have nearly identical variability levels, except for bursty-sn which shows larger fluctuations, with a maximum deviation up to 0.3-0.4 mag.
@@
We compare the results from the SPICE simulations to values extracted from observational (Ciesla et al., 2024; Endsley et al., 2024), as well as theoretical (Muñoz et al., 2023; Mason et al., 2023; Shen et al., 2023; Pallottini & Ferrara, 2023; Sun et al., 2023b, a, 2024; Kravtsov & Belokurov, 2024; Semenov et al., 2024) studies. Employing a semi-empirical approach, Muñoz et al. (2023) found a value of $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ as high as 2.5 at $z\approx 12$, which drops to 0.8 at $z\leq 10$. While none of our models reproduces such large high-$z$ values, at lower redshift there is a better agreement, although in this case our $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ are always slightly larger. Mason et al. (2023) utilized a fully empirical framework to predict that a $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ up to 1.5 in the range $108<z<$ 10 is necessary for models to align with existing observations. These numbers are similar to those predicted in our bursty-sn model. Following a similar methodology, Shen et al. (2023) found a lower value of 0.75 at $z\approx 9$, which is instead consistent with those predicted by the smooth-sn and hyper-sn models. At $z\approx 10$ and 12, though, they predict values of 1.5 and 2.0, respectively, which are higher than those of our models. Using a semi-analytical approach, Sun et al. (2024) (but see also Sun et al. 2023a, b, where bursty star formation in FIRE-2 simulation has been used to explain high redshift JWST observations) found the most probable UVLF variability value to be $\approx 1.9\pm 1.0$ at $z\approx 6$. At the same redshift, Endsley et al. (2024) found a variability in the range 1-2, obtained by analyzing Lyman-break galaxies assembled from ACS+NIRCam imaging in the GOODS and Abell 2744 fields. These values are generally reproduced in bursty-sn, as well as by halos with $M_{\rm halo}\leq\rm{10^{9}M_{\odot}}$ in the other models. Ciesla et al. (2024), using spectral energy distribution modeling with the CIGALE code, and assuming a non-parametric star formation history on the JADES public catalog, found $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle\sim 1.2$ in the range $10.56.5<z<$ 10.5, which is fairly consistent with our bursty-sn model. Pallottini & Ferrara (2023) analyzed 245 galaxies from the SERRA simulation suite at $z\approx 7.7$, finding a $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ of 0.61. We note, though, that the method employed for this evaluation differs from that used in our study and other works (see appendix A for more details). Finally, Semenov et al. (2024) investigated the importance of turbulent star formation by utilizing detailed modelling of cold turbulent ISM, star formation and feedback through zoom-in high resolution simulation of an early forming Milky Way analog, and they found that the variability decreases with time, similarly to the predictions of our models, with the exception of the $96<z<$ 9 range in bursty-sn. It is also interesting to note that the curves for the smooth-sn and hyper-sn models in the lowest mass bin aligns very well with Semenov et al. (2024) till $z\approx 7.5$, whereas after that the latter shows much lower values.

### v2

@@ Abstract
pubyear: 2015pagerange: Variability of the UV luminosity function with SPICEA
@@ 1 Introduction
[…]
@@ 2 The SPICE simulation suite
SPICE has been performed with RAMSES-RT (Rosdahl et al., 2013; Rosdahl & Teyssier, 2015), the radiation-hydrodynamics extension of the Eulerian Adaptive Mesh Refinement (AMR) code RAMSES (Teyssier, 2002). The simulations target a cubic box of size 10 $\rm{\mathit{h}^{-1}cMpc}$ with $512^{3}$ dark matter particles of mean mass $m\rm{{}_{dm}}=6.38\times 10^{5}\mathrm{M_{\odot}}$. The AMR allows each parent gas cell to split into 8 smaller cells when certain conditions are met (see B24). As a reference, this provides a physical resolution at $z=5$ ranging from 4.8 kpc for the coarsest level ($\ell_{\text{min}}=9$) to approximately 28 pc for the finest level ($\ell_{\text{max}}=16$). The hydrodynamical equations are solved using a second order Godonov scheme and the dynamics of collisionless dark matter and stellar particles are computed by solving the Poisson equation with a particle-mesh solver, projecting onto a grid using the Cloud-in-Cell scheme (Guillet & Teyssier, 2011). Gas cooling and heating are accounted for as in Rosdahl et al. (2013). Star-formation has been modelled following Kretschmer & Teyssier (2020), which adopt a subgrid turbulence model depending on the physical conditions of each computational cell including parameters like the local gas turbulent Mach number and the virial parameter, which acts as a measure of the gravitational balance in the region, and which results in a spatially-variable star formation efficiency. Additionally, SPICE includes a radiative transfer scheme with five frequency groups, i.e. infrared (0.1-1 $\rm{eV}$), optical (1-13.6 $\rm{eV}$), and three UV (13.6-24.59 $\rm{eV}$; 24.59-54.42 $\rm{eV}$; 54.42-$\infty$ $\rm{eV}$).
@@
  • "smooth-sn": for a given stellar population, a SN with mass $20\rm{M_{\odot}}$ can explode as early as 3 Myr, whereas SN with progenitor mass of $8\rm{M_{\odot}}$ can explode as late as 40 Myr. An energy of $2\times 10^{51}$ ergs per SN explosion is injected into the neighbouring cells.
@@ 3.1 UV luminosity function
Figure 1: 1500 $\mathrm{\SIUnitSymbolAngstrom}$ luminosity functions at $z$ = 12 (top panel) and 11 (bottom panel) for three feedback models (in different colors). Solid and dashed curves refer to intrinsic and dust attenuated LFs, respectively. A compilation of observations from HST and JWST (Bouwens et al., 2015; Harikane et al., 2022; Naidu et al., 2022; Adams et al., 2023; Harikane et al., 2023; Bouwens et al., 2023a, b; Leung et al., 2023; Donnan et al., 2023b, c; Pérez-González et al., 2023; Casey et al., 2024; Robertson et al., 2024; McLeod et al., 2024; Whitler et al., 2025) is shown as gray data points.
While in B24 we showed results at $z\leq 10$, in Figure 1 we present the 1500 Å LF for all feedback models at $z=11$ and 12. The LFs have been derived from the stellar spectral energy distribution within a 10 $\mathrm{\SIUnitSymbolAngstrom}$ bin centered around 1500 $\mathrm{\SIUnitSymbolAngstrom}$ . We use the SED model of BPASSv2.2.1 (Eldridge et al., 2017; Stanway & Eldridge, 2018), assuming the Chabrier (2003) IMF (details are mentioned in B24). The solid and dashed lines in the figure indicate the intrinsic and dust-attenuated luminosity functions, respectively. To account for dust attenuation, we utilize the Monte Carlo radiative transfer code RASCAS (Michel-Dansac et al., 2020). In this approach, 100 rays are cast from each stellar particle within a halo out to its virial radius. The dust attenuation along each ray is computed following the model described in the previous section. Note that, rather than applying a fixed star formation rate and UV luminosity conversion, the BPASSv2.2.1 SED model treats these two quantities independently while preserving their intrinsic correlation, thereby allowing us to capture their natural variability.
As already noted in B24, also at these redshifts smooth-sn produces a higher intrinsic luminosity function (by 0.3-0.4 dex) at all magnitudes. Conversely, the hyper-sn model yields the lowest galaxy count across all magnitude bins. The bursty-sn model produces results which are typically in between those of the other models. The effect of dust attenuation (dashed curves) starts to be visible from an absolute magnitude of $M_{\rm{UV}}\approx-15$, the value changing slightly depending on feedback model and redshift. When comparing the dust attenuated LFs to a compilation of observations from HST and JWST (Bouwens et al., 2015; Harikane et al., 2022; Naidu et al., 2022; Adams et al., 2023; Harikane et al., 2023; Bouwens et al., 2023a, b; Leung et al., 2023; Donnan et al., 2023b, c; Pérez-González et al., 2023; Casey et al., 2024; Robertson et al., 2024; McLeod et al., 2024; Whitler et al., 2025), we notice a fairly good match in the range of magnitudes covered by SPICE (except for the hyper-sn model), which, given the limited box size, does not extend to the brightest observed galaxies. For a more detailed discussion of the LF, we refer the reader to B24.
@@
To investigate the temporal evolution of the dust-attenuated UVLF on short timescales, in Figure 3 we present the UVLF in intervals of 10, 30, and 50 Myr around $z\approx 10$ and 8. While the general behaviour in the three models is qualitatively similar, important differences emerge in the nature of the UVLF variability. The UVLFs at these epochs are shaped by a combination of population growth, feedback-driven luminosity variability, dust attenuation, and the dominance of a few bright systems. At $z\approx 10$, population growth is the dominant effect, producing a consistent rise across all magnitude bins as we probe the very early stages of galaxy evolution in the simulation volume. This growth-driven behaviour is most evident in the smooth-sn and hyper-sn models. In the bursty-sn model, however, the SN feedback driven stochasticity of the star formation rate and thus of UV luminosity (we discuss this aspect later) plays comparable role, reducing the coherence of the uniform rise. Overall, at this redshift the apparent UVLF variability primarily reflects the rapid evolution of the galaxy population, rather than true short-timescale fluctuations of the UVLF. By $z\approx 8$, feedback-driven luminosity variability becomes more pronounced, with the bursty-sn model exhibiting the strongest fluctuations. In comparison, the smooth-sn model again shows a more gradual, monotonic evolution in the number of galaxies, consistent with the gentler nature of its SN feedback. The behaviour of the hyper-sn model lies in between: a combination of population growth and genuine feedback-driven variability, with the former still playing a more substantial role.
In the following sections, we investigate in more detail the origin of this feedback-driven variability.
@@
Figure 10: Redshift evolution of $\langle\sigma\rm{{}_{UV}^{dust}}\rangle$ for different halo mass bins i.e. $M\rm{{}_{halo}}>10^{10}$ M (solid lines) (top panel); $M\rm{{}_{halo}}<10^{9}$ M (dashed) and $10^{9}$ $_{\odot}<M\rm{{}_{halo}}<10^{10} M$ (dotted) (bottom). The curves for the entire sample are shown in the top panel as solid lines with black edges. Colors refer to the three SN feedback models, while the values from observational and other theoretical studies (Pallottini & Ferrara, 2023; Shen et al., 2023; Ciesla et al., 2024; Muñoz et al., 2023; Mason et al., 2023; Kravtsov & Belokurov, 2024; Sun et al., 2024; Endsley et al., 2024; Semenov et al., 2024) are shown in grey data points and shaded regions in both the panels.
@@ 3.2 Star formation rate variability
For a more quantitative investigation of the SFR variability, we first compute the median star-formation main sequence in different halo mass bins and redshift intervals (corresponding to bins of 100 Myr [^1] This has been calculated from the SFR values of all the halos present in each 100 Myr time intervals and then computing the median SFR as a function of halo mass. We have also performed the same analysis using bin sizes of 30 and 50 Myr, and found no qualitative differences in our conclusions.), SFRmedian. In the inset of Figure 4 we show the variability of the SFR for the same halo, defined as $\rm{\mathit{\delta}_{SFR}}={\rm log}_{10}\rm{(SFR/SFR_{median})}$, which quantifies how the SFR of the halo deviates from the median SFR computed over the entire sample of halos in the same mass bin. In the inset of Figure 4 we show the distribution of $\delta_{\rm SFR}$ for this halo during its entire lifetime, together with the corresponding standard deviation, $\sigma_{\rm SFR}$. As expected, bursty-sn shows the widest distribution, with $\sigma_{\rm SFR,bu}=0.34$, which is more than double the one of smooth-sn with $\sigma_{\rm SFR,sm}=0.12$. hyper-sn produces a $\delta_{\rm SFR}$ distribution which is similar to the one from smooth-sn, with a slightly higher standard deviation of $\sigma_{\rm SFR,hn}=0.14$, reflecting the stronger fluctuations present at $z>9$.  [^2] This behaviour persists when comparing objects of similar stellar and halo masses across the models with bursty-sn continuing to produce the largest scatter. The level of variability could, in principle, depend on the size of the simulated volume, decreasing as fluctuations from individual objects partially cancel out in larger boxes. To test that, we repeated the analysis on smaller subvolumes (i.e. by selecting objects within cropped regions of the full box) and found out the variability remains essentially unchanged across different portions of the simulation, indicating that the box-size effect is minimal in this case. Larger boxes could still include higher overdensities and physical processes such as black-hole feedback, which might change the variability, but these are not included in the current SPICE simulations.
@@ 3.3 Variability of the UV luminosity function
Figure 7 shows the scatter in UV luminosity, $\sigma_{\rm UV}^{\rm dust}$, as a function of $M_{\rm halo}$ in the range z=515z=5-15. At the highest redshifts, all models exhibit a similar dependence on halo mass, with smaller halos showing a larger UVLF variability, likely due to efficient SN feedback in this mass regime (Gelli et al., 2024). As expected, bursty-sn has the highest variability among all models, with a UV luminosity scatter which starts to rise at $z\lesssim 10$ and reaches $\sigma_{\rm UV}^{\rm dust}\approx 2.5$ for $M_{\rm halo}\approx 10^{10.3}$ M at $z\approx 6.5$. At lower redshift, $\sigma_{\rm UV}^{\rm dust}$ decreases again. In comparison, for smooth-sn the UV luminosity scatter is always below $\sigma_{\rm UV}^{\rm dust}\approx 1.3$, whereas hyper-sn shows slightly higher variability, with a maximum value of $\sigma_{\rm UV}^{\rm dust}\approx 1.5$. Indeed, hyper-sn exhibits a variability comparable to the one in bursty-sn until $z\approx 9$, while at lower redshift the variability decreases as a result of the reduced fraction of HN, and becomes more similar to the smooth-sn one. The same behaviour had already been observed in Figure 4 with respect to the SFR. At $z<6$ all models have nearly identical variability levels, except for bursty-sn which shows larger fluctuations, with a maximum deviation up to 0.3-0.4 mag.
@@
We compare the results from the SPICE simulations to values extracted from observational (Ciesla et al., 2024; Endsley et al., 2024), as well as theoretical (Muñoz et al., 2023; Mason et al., 2023; Shen et al., 2023; Pallottini & Ferrara, 2023; Sun et al., 2023b, a, 2024; Kravtsov & Belokurov, 2024; Semenov et al., 2024) studies. Employing a semi-empirical approach, Muñoz et al. (2023) found a value of $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ as high as 2.5 at $z\approx 12$, which drops to 0.8 at $z\leq 10$. While none of our models reproduces such large high-$z$ values, at lower redshift there is a better agreement, although in this case our $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ are always slightly larger. Mason et al. (2023) utilized a fully empirical framework to predict that a $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ up to 1.5 in the range $8<z<10$ is necessary for models to align with existing observations. These numbers are similar to those predicted in our bursty-sn model. Following a similar methodology, Shen et al. (2023) found a lower value of 0.75 at $z\approx 9$, which is instead consistent with those predicted by the smooth-sn and hyper-sn models. At $z\approx 10$ and 12, though, they predict values of 1.5 and 2.0, respectively, which are higher than those of our models. Using a semi-analytical approach, Sun et al. (2024) (but see also Sun et al. 2023a, b, where bursty star formation in FIRE-2 simulation has been used to explain high redshift JWST observations) found the most probable UVLF variability value to be $\approx 1.9\pm 1.0$ at $z\approx 6$. At the same redshift, Endsley et al. (2024) found a variability in the range 1-2, obtained by analyzing Lyman-break galaxies assembled from ACS+NIRCam imaging in the GOODS and Abell 2744 fields. These values are generally reproduced in bursty-sn, as well as by halos with $M_{\rm halo}\leq\rm{10^{9}M_{\odot}}$ in the other models. Ciesla et al. (2024), using spectral energy distribution modeling with the CIGALE code, and assuming a non-parametric star formation history on the JADES public catalog, found $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle\sim 1.2$ in the range $6.5<z<10.5$, which is fairly consistent with our bursty-sn model. Pallottini & Ferrara (2023) analyzed 245 galaxies from the SERRA simulation suite at $z\approx 7.7$, finding a $\langle\sigma\rm{{}_{UV}^{\rm dust}}\rangle$ of 0.61. We note, though, that the method employed for this evaluation differs from that used in our study and other works (see appendix A for more details). Finally, Semenov et al. (2024) investigated the importance of turbulent star formation by utilizing detailed modelling of cold turbulent ISM, star formation and feedback through zoom-in high resolution simulation of an early forming Milky Way analog, and they found that the variability decreases with time, similarly to the predictions of our models, with the exception of the $6<z<9$ range in bursty-sn. It is also interesting to note that the curves for the smooth-sn and hyper-sn models in the lowest mass bin aligns very well with Semenov et al. (2024) till $z\approx 7.5$, whereas after that the latter shows much lower values.